A long, straight wire carries a current of 5.20 A. An electron is traveling in the vicinity of the wire.At the instant when the electron is 4.70 cm from the wire and traveling with a speed of 5.80×10^4 m/s directly toward the wire, what is the magnitude of the force that the magnetic field of the current exerts on the electron?
Answer #1

The magnitic field due to an infinite long current carrying conductor is B= \frac{\mu _{0}i}{2\pi r}

From Biot-Savart law \bar{dB}= \frac{\mu _{0}i\bar{dl}x\hat{r}}{2\pi r}

Therefore     B= \frac{4\pi x10^{-7}xi}{2\pi r}

                    B= 2.21x10-5T

Now the force on moving charge is F=QVB

           Force on electron F= 1.6x10-19x5.8x104 x2.21x10-5

                                      F = 20.51x10-20N

It is perpedicular to the conductor.